OpenStudy (anonymous):

please help me!?!!!?! What is the solution of the equation? (equation below in a comment)

OpenStudy (anonymous):

\[\sqrt{2x+13}-5=x\]

OpenStudy (anonymous):

one is -2

OpenStudy (anonymous):

how did u get that tho?

OpenStudy (anonymous):

take -5 to other side of = and square both sides

OpenStudy (anonymous):

fio dont get it can u show me in an equation please im very lost here.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

is x=-6, -2 right?